OpenStudy (anonymous):

What is the diameter of a cone with height 8 m and volume 150pi m^3?

OpenStudy (bahrom7893):

V = (1/3)pi*r^2*h 150=(1/3)pi*(d/2)^2*8 because r = d/2 (radius is half the diameter)

OpenStudy (bahrom7893):

Solve for d and you're done.

OpenStudy (anonymous):

Is it 7.5 sqrt3? @bahrom7893

OpenStudy (bahrom7893):

Could you post the steps please?

OpenStudy (anonymous):

Nvm I got it

OpenStudy (bahrom7893):

Great